How much will the wire stretch? Help? How much will the wire stretch? Help?

How much will the wire stretch? Help?
Answer #1

Here , for aluminium

Young Modulus , Y = 6.9 *10^13 Pa

Now , let the wire streching length is l

as Y = stress/strain

strain = stress/Y

l/L = F/A*Y

l/2.5 = 4*9.8/(6.9 *10^13 * 3.141 * (0.0015/2)^2)

solving for l

l = 2.545 *10^-6 m

the wire strech by 2.545 *10^-6 m